# UniVideo: Unified Understanding, Generation, and Editing for Videos

### v2

@@ 3.3.1 Zero Shot Generalization
We observed two type of generalization ability of UniVideo. Although the training data of UniVideo does not include general free-form video editing tasks (see Table 1), it transfers this ability from diverse image editing data and in-context video editing data (limited to ID deletion, swapping, addition, and stylization) to the video domain, enabling it to handle free-form video editing instructions(e.g., changing material or environment). Surprisingly, we find that UniVideo can perform tasks such as green-screening characters from videos. We also observe that UniVideo is capable of handling task compositions. It can combine in-context editing with style transfer, or perform multiple edits simultaneously (e.g., deleting one identity while adding another). Demonstrations in Figure 5.
@@ 3.4 Ablation Study
Our ablation studies address two central questions: (i) Does multi-task learning enhance performance compared with single-task learning? (ii) Is our model design effective? Specifically, should visual embeddings be streamed to both the MLLM and MMDiT branches? We conduct human evaluations on In-Context Video Editing and In-Context Video Generation, using the same evaluation protocol as in subsubsection 3.2.2. (i) To study multi-task learning, we compare UniVideo with a single-task baseline. The single-task baseline shares the same architecture as UniVideo but requires an independent model for each task and has access only to task-specific data. Results in Table 6 demonstrate the effectiveness of multi-task learning, especially for the editing task, where UniVideo benefits from large-scale image editing data during joint learning. (ii) To evaluate the impact of streaming visual inputs, we compare UniVideo with a variant that share the same architecture: - w/o visual for MMDiT: visual inputs are fed only to the MLLM branch. As shown in Table 7, feeding visual inputs exclusively to the MLLM results in a dramatic drop in identity preservation.
@@ 4 Related Work
Image/Video Generation and Editing. Diffusion models have achieved remarkable success in high-fidelity image synthesis, with systems like Stable Diffusion, DALL·E, and Imagen(Rombach et al., 2022; Podell et al., 2023; Esser et al., 2024; Ramesh et al., 2021; Saharia et al., 2022) establishing strong text-to-image capabilities and recent video diffusion models(Blattmann et al., 2023b; Polyak et al., 2024; Chen et al., 2025c; 2023; Yang et al., 2024; Blattmann et al., 2023a; Kong et al., 2024; Brooks et al., 2024; Ma et al., 2025a) enabling scalable video generation. To improve controllability, models including ControlNet, T2I-Adapter(Zhang et al., 2023b; Mou et al., 2024) introduce external condition modules, while editing frameworks like InstructPix2Pix, EMU-Edit (Brooks et al., 2023; Sheynin et al., 2024) support instruction-driven refinement. Recently, unified image generation has emerged, with OmniGen, OmniControl, and UniReal (Xiao et al., 2025; Tan et al., 2024; Chen et al., 2025d) expanding from generation to reference-guided editing. General editing methods (Wei et al., 2024; Zhao et al., 2024; Liu et al., 2025b; Shi et al., 2024b; Zhang et al., 2023a) further highlight this trend. In contrast, the video domain remains dominated by single-task frameworks such as Video-P2P, MagicEdit, MotionCtrl (Liu et al., 2024c; Liew et al., 2023; Wang et al., 2024c; Liu et al., 2025a). Attempts at unification include AnyV2V(Ku et al., 2024), which requires task-specific pipelines, VACE(Jiang et al., 2025), which relies on heavy adapter designs, FullDiT(Ju et al., 2025), which supports multi-condition video generation but lacks editing, and UNIC(Ye et al., 2025b), which unifies tasks but depends on task-specific condition bias, limiting scalability. Yet, compared to images, unified and flexible video generation and editing remains far less explored. Our work bridges this gap by unifying diverse video tasks under a multimodal instruction framework. We provide the model capabilities comparison in Table 8.
@@ Appendix B Training Details
We adopt qwen2.5VL-7B (Bai et al., 2025) as the MLLM backbone and HunyuanVideo-T2V-13B (Kong et al., 2024) as the MMDiT backbone. The original HunyuanVideo also uses CLIP as its text encoder; we remove it and instead employ qwen2.5VL as the unified multimodal embedder. The released HunyuanVideo checkpoint is a CFG-distilled model, whose distillation embeddings we discard to simplify the training. To align feature dimensions between qwen2.5VL and HunyuanVideo, we apply an MLP with a $4\times$ expansion. Training is conducted on 32 H100 GPUs. We report training configurations, hyperparameters, and data composition ratios in Table 2, and provide task example quantity in Table 1.
@@ D.3 Image and Video
We leverage state-of-the-art image editing models such as FLUX.1 Kontext (Labs et al., 2025) to create diverse image editing data. We also source open source data such as OmniEdit(Wei et al., 2024), ImgEdit(Ye et al., 2025a), and ShareGPT-4o-Image(Chen et al., 2025b). For image and video generation tasks, we utilize internal datasets. The detailed data composition is provided in Table 1.

### v3

@@ 3.4.1 Zero Shot Generalization
We observed two type of generalization ability of UniVideo. Although the training data of UniVideo does not include general free-form video editing tasks, it transfers this ability from diverse image editing data and in-context video editing data (limited to ID deletion, swapping, addition, and stylization) to the video domain, enabling it to handle free-form video editing instructions(e.g., changing material or environment). Surprisingly, we find that UniVideo can perform tasks such as changing materials of character. We also observe that UniVideo is capable of handling task compositions. It can combine in-context editing with style transfer, or perform multiple edits simultaneously (e.g., deleting one identity while adding another). Demonstrations in Figure 8.
@@ 3.5 Ablation Study
Our ablation studies address the following two questions: (i) Does multi-task learning enhance performance compared with single-task learning? (ii) Is our two-branch design effective? Specifically, should visual embeddings be streamed to both the MLLM and MMDiT branches?
We conduct human evaluations on In-Context Video Editing and In-Context Video Generation, using the same evaluation protocol as in subsubsection 3.3.2. (i) To study multi-task learning, we compare UniVideo with a single-task baseline. The single-task baseline shares the same architecture as UniVideo but requires an independent model for each task and has access only to task-specific data. Results in Table 5 demonstrate the effectiveness of multi-task learning, especially for the editing task, where UniVideo benefits from large-scale image editing data during joint learning. (ii) To evaluate the impact of streaming visual inputs, we compare UniVideo with a variant that share the same architecture: - w/o visual for MMDiT: visual inputs are fed only to the MLLM branch. As shown in Table 6, feeding visual inputs exclusively to the MLLM results in a dramatic drop in identity preservation.
@@ 4 Related Work
Image/Video Generation and Editing. Diffusion models have achieved remarkable success in high-fidelity image synthesis, with systems like Stable Diffusion, DALL·E, and Imagen(Rombach et al., 2022; Podell et al., 2023; Esser et al., 2024; Ramesh et al., 2021; Saharia et al., 2022) establishing strong text-to-image capabilities and recent video diffusion models(Blattmann et al., 2023b; Polyak et al., 2024; Chen et al., 2025c; 2023; Yang et al., 2024; Blattmann et al., 2023a; Kong et al., 2024; Brooks et al., 2024; Ma et al., 2025a) enabling scalable video generation. To improve controllability, models including ControlNet, T2I-Adapter(Zhang et al., 2023b; Mou et al., 2024) introduce external condition modules, while editing frameworks like InstructPix2Pix, EMU-Edit (Brooks et al., 2023; Sheynin et al., 2024) support instruction-driven refinement. Recently, unified image generation has emerged, with OmniGen, OmniControl, and UniReal (Xiao et al., 2025; Tan et al., 2024; Chen et al., 2025d) expanding from generation to reference-guided editing. General editing methods (Wei et al., 2024; Zhao et al., 2024; Liu et al., 2025b; Shi et al., 2024b; Zhang et al., 2023a) further highlight this trend. In contrast, the video domain remains dominated by single-task frameworks such as Video-P2P, MagicEdit, MotionCtrl (Liu et al., 2024c; Liew et al., 2023; Wang et al., 2024c; Liu et al., 2025a). Attempts at unification include AnyV2V(Ku et al., 2024), which requires task-specific pipelines, EditVerse(Ju et al., 2025a), which can not perform visual understanding task. VACE(Jiang et al., 2025), which relies on heavy adapter designs, FullDiT(Ju et al., 2025b), which supports multi-condition video generation but lacks editing, and UNIC(Ye et al., 2025b), which unifies tasks but depends on task-specific condition bias, limiting scalability. Yet, compared to images, unified and flexible video generation and editing remains far less explored. Our work bridges this gap by unifying diverse video tasks under a multimodal instruction framework. We provide the model capabilities comparison in Table 7.
@@ Appendix B Training Details
We adopt qwen2.5VL-7B (Bai et al., 2025) as the MLLM backbone and HunyuanVideo-T2V-13B (Kong et al., 2024) as the MMDiT backbone. The original HunyuanVideo also uses CLIP as its text encoder; we remove it and instead employ qwen2.5VL as the unified multimodal embedder. The released HunyuanVideo checkpoint is a CFG-distilled model, whose distillation embeddings we discard to simplify the training. To align feature dimensions between qwen2.5VL and HunyuanVideo, we apply an MLP with a $4\times$ expansion. We report training configurations, hyperparameters in Table 1
@@ D.3 Image and Video
We leverage state-of-the-art image editing models such as FLUX.1 Kontext (Labs et al., 2025) to create diverse image editing data. We also source open source data such as OmniEdit(Wei et al., 2024), ImgEdit(Ye et al., 2025a), and ShareGPT-4o-Image(Chen et al., 2025b). For image and video generation tasks, we utilize internal datasets.